# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

### v2

@@ 3.2 A priori estimates and uniqueness
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\mathbb{E}\left[\left|K_{T}\right|^{2}\right]\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}\left(S_{t}\right)^{+}\right|^{2}\right).\end{split}$
@@
$\begin{split}&\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\theta\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mu\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &=\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}f(s,Y_{s},Z_{s})ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}h(s,Y_{s})d\kappa_{s}\\ &\quad+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds-2\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s-}Z^{(k)}_{s}dH^{(k)}_{s}\\ &\quad-\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}Z_{s}^{(k)}Z_{s}^{(k^{\prime})}\,\left(d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}-d\big\langle H^{(k)},H^{(k^{\prime})}\big\rangle_{s}\right)+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}dK_{s}.\end{split}$ (3.10)
@@
$\begin{split}2Y_{s}f(s,Y_{s},Z_{s})ds&\leq 2\lambda_{s}\left|Y_{s}\right|^{2}ds+2\left|Y_{s}\right|\left(\varphi_{s}+\eta_{s}\left\|\gamma_{s}Z_{s}\right\|_{\ell^{2}}\right)ds\\ &\leq\left\{2\lambda_{s}+\frac{1}{\varepsilon^{\prime}}\eta_{s}^{2}+\varepsilon a^{2}_{s}\right\}\left|Y_{s}\right|^{2}ds+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\left(2+\varepsilon+\frac{1}{\varepsilon^{\prime}}\right)\left|Y_{s}\right|^{2}dV_{s}+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds,\end{split}$
and
$\begin{split}2Y_{s}h(s,Y_{s})d\kappa_{s}&\leq 2\varrho_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+2\left|Y_{s}\right|\psi_{s}d\kappa_{s}\\ &\leq\varepsilon^{\prime\prime}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{1}{\varepsilon^{\prime\prime}}\varphi_{s}^{2}d\kappa_{s},\end{split}$
and
$\begin{split}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\left(\rho_{s}\left|Y_{s}\right|^{2}+\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\left|Y_{s}\right|^{2}dV_{s}+2\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds+2\left|g(s,0,0)\right|^{2}ds.\end{split}$
@@
$\begin{split}&\mathbb{E}\left[\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}\right]+\left(\theta-4-\varepsilon-\frac{1}{\varepsilon^{\prime}}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}\\ &\qquad+\left(\mu-\varepsilon^{\prime\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+(1-2\alpha-\varepsilon^{\prime})\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathbb{E}\left[\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}\right]+\frac{1}{\varepsilon}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\frac{1}{\varepsilon^{\prime\prime}}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\\ &\qquad+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}S_{s}dK_{s}.\end{split}$
@@
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\right)\\ &\qquad+\varepsilon^{\prime\prime\prime}\mathfrak{c}_{\theta,\mu,\alpha}^{2}\mathbb{E}\sup_{t\leq u\leq T}\left|\Phi^{\theta,\mu}_{u}(S_{u})^{+}\right|^{2}+\frac{1}{\varepsilon^{\prime\prime\prime}}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}.\end{split}$ (3.11)
@@
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\left(\mathbb{E}\left|Y_{t}\right|^{2}+\mathbb{E}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\left.+\mathbb{E}\left(\int_{t}^{t}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}\right)^{2}+\mathbb{E}\left(\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s}\right)^{2}\right)\\ &\leq 6\left(\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\qquad\left.+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right).\end{split}$ (3.12)
@@
$\begin{split}\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{\theta,-\mu}_{s}d\kappa_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|h(s,Y_{s})\right|^{2}d\kappa_{s}\right)\\ &\leq\frac{2\zeta^{2}}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{2}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\psi^{2}_{s}d\kappa_{s}.\end{split}$
@@
$\begin{split}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(\left|Y_{s}\right|^{2}dV_{s}+\alpha\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds.\end{split}$
@@
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\mathbb{E}\left|Y_{t}\right|^{2}+6\mathbb{E}\left|\xi\right|^{2}+\left(\frac{18}{\theta}+\frac{12\zeta^{2}}{\mu}+12\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}\\ &\qquad+\left(\frac{18}{\theta}+6\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\frac{18}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds\\ &\qquad+12\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds.\end{split}$ (3.13)
@@
$\begin{split}&\sup_{0\leq t\leq T}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$ (3.14)
@@
$\begin{split}\mathbb{E}\left|K_{T}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$
Now, applying the Burkholder-Davis–Gundy inequality similar to those used in Corollary 7, we have
@@
Finally, the Burkholder-Davis-Gundy inequality along with the Kunita-Watanabe inequality (see, e.g., [42, Ch II. Theorem 25]), we get
@@
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right).\end{split}$

### v3

@@ 3.2 A priori estimates and uniqueness
$\begin{split}&\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\mathbb{E}\left[\left|K_{T}\right|^{2}\right]\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}\left(S_{t}\right)^{+}\right|^{2}\right)\end{split}$
@@
$\begin{split}&\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\theta\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mu\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &=\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}f(s,Y_{s},Z_{s})ds+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}h(s,Y_{s})d\kappa_{s}\\ &\quad+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds-2\sum_{k=1}^{d}\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s-}Z^{(k)}_{s}dH^{(k)}_{s}\\ &\quad-\sum_{k,k^{\prime}=1}^{d}\int_{0}^{t}\Phi^{\theta,\mu}_{s}Z_{s}^{(k)}Z_{s}^{(k^{\prime})}\,\left(d\big[H^{(k)},H^{(k^{\prime})}\big]_{s}-d\big\langle H^{(k)},H^{(k^{\prime})}\big\rangle_{s}\right)+2\int_{t}^{T}\Phi^{\theta,\mu}_{s}Y_{s}dK_{s}\end{split}$ (3.10)
@@
$\begin{split}2Y_{s}f(s,Y_{s},Z_{s})ds&\leq 2\lambda_{s}\left|Y_{s}\right|^{2}ds+2\left|Y_{s}\right|\left(\varphi_{s}+\eta_{s}\left\|\gamma_{s}Z_{s}\right\|_{\ell^{2}}\right)ds\\ &\leq\left\{2\lambda_{s}+\frac{1}{\varepsilon^{\prime}}\eta_{s}^{2}+\varepsilon a^{2}_{s}\right\}\left|Y_{s}\right|^{2}ds+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\left(2+\varepsilon+\frac{1}{\varepsilon^{\prime}}\right)\left|Y_{s}\right|^{2}dV_{s}+\frac{1}{\varepsilon}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\varepsilon^{\prime}\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\end{split}$
$\begin{split}2Y_{s}h(s,Y_{s})d\kappa_{s}&\leq 2\varrho_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+2\left|Y_{s}\right|\psi_{s}d\kappa_{s}\\ &\leq\varepsilon^{\prime\prime}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{1}{\varepsilon^{\prime\prime}}\varphi_{s}^{2}d\kappa_{s}\end{split}$
$\begin{split}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\left(\rho_{s}\left|Y_{s}\right|^{2}+\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\left|Y_{s}\right|^{2}dV_{s}+2\alpha\left\|\gamma_{s}Z_{s}\right\|^{2}_{\ell^{2}}ds+2\left|g(s,0,0)\right|^{2}ds\end{split}$
@@
$\begin{split}&\mathbb{E}\left[\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}\right]+\left(\theta-4-\varepsilon-\frac{1}{\varepsilon^{\prime}}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}\\ &\qquad+\left(\mu-\varepsilon^{\prime\prime}\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+(1-2\alpha-\varepsilon^{\prime})\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathbb{E}\left[\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}\right]+\frac{1}{\varepsilon}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\frac{1}{\varepsilon^{\prime\prime}}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\\ &\qquad+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}S_{s}dK_{s}\end{split}$
@@
$\begin{split}&\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\right)\\ &\qquad+\varepsilon^{\prime\prime\prime}\mathfrak{c}_{\theta,\mu,\alpha}^{2}\mathbb{E}\sup_{t\leq u\leq T}\left|\Phi^{\theta,\mu}_{u}(S_{u})^{+}\right|^{2}+\frac{1}{\varepsilon^{\prime\prime\prime}}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}\end{split}$ (3.11)
@@
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\left(\mathbb{E}\left|Y_{t}\right|^{2}+\mathbb{E}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\left.+\mathbb{E}\left(\int_{t}^{t}g(s,Y_{s},Z_{s})\overleftarrow{dB}_{s}\right)^{2}+\mathbb{E}\left(\sum_{k=1}^{d}\int_{t}^{T}Z^{(k)}_{s}dH^{(k)}_{s}\right)^{2}\right)\\ &\leq 6\left(\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\left(\int_{t}^{T}f(s,Y_{s},Z_{s})ds\right)^{2}+\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}\right.\\ &\qquad\left.+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right)\end{split}$ (3.12)
@@
$\begin{split}\mathbb{E}\left(\int_{t}^{T}h(s,Y_{s})d\kappa_{s}\right)^{2}&\leq\mathbb{E}\left(\int_{t}^{T}\Phi^{\theta,-\mu}_{s}d\kappa_{s}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|h(s,Y_{s})\right|^{2}d\kappa_{s}\right)\\ &\leq\frac{2\zeta^{2}}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\kappa_{s}+\frac{2}{\mu}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\psi^{2}_{s}d\kappa_{s}\end{split}$
@@
$\begin{split}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,Y_{s},Z_{s})\right|^{2}ds&\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left(\left|Y_{s}\right|^{2}dV_{s}+\alpha\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\right)+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\\ &\leq 2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}dV_{s}+\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+2\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\end{split}$
@@
$\begin{split}\mathbb{E}\left|K_{T}-K_{t}\right|^{2}&\leq 6\mathbb{E}\left|Y_{t}\right|^{2}+6\mathbb{E}\left|\xi\right|^{2}+\left(\frac{18}{\theta}+\frac{12\zeta^{2}}{\mu}+12\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}\\ &\qquad+\left(\frac{18}{\theta}+6\right)\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds+\frac{18}{\theta}\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds\\ &\qquad+12\mathbb{E}\int_{t}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds\end{split}$ (3.13)
@@
$\begin{split}&\sup_{0\leq t\leq T}\mathbb{E}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|Y_{s}\right|^{2}d\mathcal{Q}_{s}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left\|Z_{s}\gamma_{s}\right\|^{2}_{\ell^{2}}ds\\ &\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right)\end{split}$ (3.14)
@@
$\begin{split}\mathbb{E}\left|K_{T}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right)\end{split}$
Now, applying the B-D–G inequality similar to those used in Corollary 8, we have
@@
Finally, the B-D-G inequality along with Moreover, using again the B-D-G inequality along with the Kunita-Watanabe inequality (see, e.g., [41, Ch II. Theorem 25]), we get
@@
$\begin{split}\mathbb{E}\sup_{0\leq t\leq T}\Phi^{\theta,\mu}_{t}\left|Y_{t}\right|^{2}&\leq\mathfrak{c}_{\theta,\mu,\alpha,\zeta}\left(\mathbb{E}\Phi^{\theta,\mu}_{T}\left|\xi\right|^{2}+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\frac{\varphi_{s}}{a_{s}}\right|^{2}ds+\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|\psi_{s}\right|^{2}d\kappa_{s}\right.\\ &\qquad\left.+\mathbb{E}\int_{0}^{T}\Phi^{\theta,\mu}_{s}\left|g(s,0,0)\right|^{2}ds+\mathbb{E}\sup_{0\leq t\leq T}\left|\Phi^{\theta,\mu}_{t}(S_{t})^{+}\right|^{2}\right)\end{split}$